OpenStudy (anonymous):

Please Help Find the first six terms of the sequence. a1 = -8, an = 5 • a(n-1)

OpenStudy (anonymous):

n=2, you have an is a2, right? and a(n-1) is a1, right? so a2 = 5 a1 = 5*(-8) =-40 what if n =3?? do the same, show me your work

OpenStudy (anonymous):

5*-8(3-1) 5*-8(2) 5*-16 -80

OpenStudy (anonymous):

-8, -40, -200, -1000, -5000, -25,000 -8, -40, -35, -30, -25, -20 0, 5, -40, -35, -30, -25 -40, -200, -1000, -5000, -25,000, -125,000

OpenStudy (anonymous):

but -80 isn't one of the answers

OpenStudy (anonymous):

@Bebong

OpenStudy (anonymous):

of course it is not. hihihi. you apply the wrong rule.

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

\(a_n= 5*a_{n-1}\) , it is not \(a_n = 5 *(n-1)\)

OpenStudy (anonymous):

so would a3 = -200

OpenStudy (anonymous):

if n =3, then \(a_n ~~is~~a_3\) and \(a_{n-1}~~is~~a_2\) you have a2 from the previous calculation, right? plug in

OpenStudy (anonymous):

let check, a3 = 5*a2 =5*(-40) =-200 , yes!! you are correct

OpenStudy (anonymous):

Thanks, I got it right!!